OpenStudy (anonymous):

graph the function by hand. give the coordinates of the vertex. f(x)= -3(x-3)^2

OpenStudy (campbell_st):

the parabola is concave down with a vertex at (3, 0)

OpenStudy (anonymous):

how did you find that answer?

OpenStudy (campbell_st):

and y intercept is y = -27

OpenStudy (campbell_st):

there are lots of ways in the factorised form given y = (x -h)^2 + k the vertex is (h, k) in your question h = 3 and k = 0 in the parabola y = ax^2 + bx + x a determines concavity a : positive concave up a negative concave down in your question a = -3 which comes from y = -3(x^2 -6x + 9) so concave down

OpenStudy (campbell_st):

y -intercept.... let x = 0